OpenStudy (anonymous):

HELP WILL MEDAL! -x+y-2z=-5 5x+4y-4z=4 x-3y-2z=3 so far i have y=1 I know the answer is x=-4 y=1 z=-5 but I don't know where to go from where I am.

OpenStudy (anonymous):

x-3y-2z=3 -x+ y+2z=-5 -2y =-2 -2 =-2 Divided by -2 y=1

OpenStudy (anonymous):

replace \(y\) by \(1\) in any two of the three equations then you will have two variables with two equations to solve

OpenStudy (texaschic101):

take 2 of the 3 equations, it does not matter which 2, and eliminate a variable -x + y - 2z = -5 x - 3y - 2z = 3 -----------------add -2y - 4z = -2 now take 2 equations, make sure you get the equation you have not messed with yet, and eliminate THE SAME variable -x + y - 2z = -5 --->(5) 5x + 4y - 4z = 4 --------------- -5x + 5y - 10z = -25 (result of multiplying by 5) 5x + 4y - 4z = 4 ---------------add 9y - 14z = - 21 now take your 2 answers and eliminate a variable -2y - 4z = -2 --->(9) 9y - 14z = -21 --->(2) --------------- -18y - 36z = -18 (result of multiplying by 9) 18y - 28z = - 42 (result of multiplying by 2) --------------add - 64z = - 60 z = -60/-64 z = 15/16 now enter 15/16 in for z in an equation that you eliminated the x out of -2y - 4z = -2 -2y - 4(15/16) = -2 -2y - 60/16 = -2 -2y = 60/16 - 2 ---> common denominator is 16 -2y = 60/16 - 32/16 -2y = 28/16 y = (28/16) / -2 y = 28/16 * -1/2 y = - 28/32 y = - 7/8 now sub 15/16 in for z and -7/8 in for y in one of the original equations -x + y - 2z = -5 -x + (-7/8) - 2(15/16) = -5 -x - 7/8 - 30/16 = -5 -x - 14/16 - 30/16 = -5 -x - 44/16 = -5 -x = 44/16 - 5 --> common denominator is 16 -x = 44/16 - 80/16 -x = - 36/16 x = 36/16 x = 9/4 you can then check your answers by subbing in to one of your original equations x - 3y - 2z = 3 9/4 - 3(-7/8) - 2(15/16) = 3 9/4 + 21/8 - 30/16 = 3 ---> common denominator is 16 36/16 + 42/16 - 30/16 = 3 78/16 - 3016 = 3 48/16 = 3 3 = 3 (correct) so x = 9/4, y = -7/8, and z = 15/16 I do not know where you got your answers from....unless the problem is written incorrectly

OpenStudy (anonymous):

x should equal -4 y should equal 1 z should equal -5

OpenStudy (texaschic101):

lets check that... if x = -4 and y = 1 and z = -5 -x + y - 2z = -5 -(-4) + 1 - 2(-5) = -5 4 + 1 + 10 = -5 15 = -5 (this is incorrect) Those are not the answers to your problem

OpenStudy (anonymous):

x-3y-2z=3 -x+y+2z=-5 ______________ -2y = -2 divide by -2 y=1

OpenStudy (anonymous):

5(-x+y+2z=-5) 5x+4y-4z=4 ________________ -5x+5y+10z=-25 5x+4y-4z=4 _________________ 9y+6z=-2 9(1)+6z=-21 9+6z=-21 -9 =-9 6z=-30 Divide by 6 z=-5

OpenStudy (anonymous):

-x+y+2z=-5 -x+(1)+2(-5)=-5 -x+1-10=-5 -x-9=-5 +9=+9 -x = 4 -*1 = *-1 x = -4

OpenStudy (texaschic101):

you wrote the problem wrong...in the original way it is posted, you have all 3 z's as being negative